OpenStudy (anonymous):

A charge q is divided into two parts q1 and (q-q1) .What is the ratio of q/q1 so that the force between the two parts placed at a given distance apart is maximum. options are - a)1:1 b) 2:1 c)1:2 d)1:4

OpenStudy (anonymous):

the way i did it was - F = kq1(q-q1)/ r^2; this force has to be maximum ; so differentiation of maxima = 0; therefore differentiation of F = 0; but i have no clue how to differentiate it..:(

OpenStudy (anonymous):

@mos1635 - help me again please..!

OpenStudy (anonymous):

The answer is 1:1 \[dF/dq1= 0\] for maximum force. \[dF/dq1= (k/r ^{2} ) * (q-2q1) = 0\] Therefore \[q-2q1=0\] \[q _{1}\] Therefore both charges are q/2

OpenStudy (anonymous):

q1=q/2

OpenStudy (mos1635):

F=K/r^2*q1(q-q1) F=K/r^2*(q*q1-q1^2) dF/dq1=k/r^2*[d/dq1(q*q1-q1^2)] dF/dq1=k/r^2*[d/dq1(q*q1)-d/dq1q1^2)] dF/dq1=k/r^2*(q-2q1) to be 0 q-2q1=0 q/q1=2

OpenStudy (anonymous):

Sorry the answer is q/q1=2 . I made a mistake in first line of my solution. I misread the question

OpenStudy (anonymous):

thanks guys..:D